Show your work CLEARLY and COMPLETELY! For z-scores: Round to the SECOND decimal place. For Probabilities: Keep the Original FOUR decimal places. For EACH part of the question: Draw a Normal Curve and highlight the corresponding area(s)! Name: Problem 1 Stanford-Binet lQ Test scores are normally distributed with a mean score of 100 and a standard deviation of 16. Find the probability that a randomly selected person has an IQ test score a) Over 130. Answer Canvas Quiz 5-1 and 5-2. (2 points) b) Under 85. Answer Canvas Quiz 5-3 and 5-4. (2 points) c) Between 72 and 128. Answer Canvas Quiz 5-5 to 5-7. (3 points) d) Between+1.25 standard deviations of the mean. Answer Canvas Quiz 5-8. (1 point)

Problem 2 A tire company has developed a new type of steel-belted radial tire. Extensive testing indicates the population of mileages obtained by all tires of this new type is normally distributed with a mean of 40,000 miles and a standard deviation of 4,000 miles. The company wishes to offer a guarantee providing a discount on a new set of tires if the original tires purchased qualify the guarantee mileage. What should the guaranteed mileage be if the tire company desires that no more than 3 percent of the tires will fail to meet the guaranteed mileage? Answer Canvas Quiz 5-9 and 5-10. (2 points)
Answer #1

a)

for normal distribution z score =(X-μ)/σ
here mean=       μ= 100
std deviation   =σ= 16.0000
probability = P(X>130) = P(Z>1.88)= 1-P(Z<1.88)= 1-0.9699= 0.0301

b)

probability = P(X<85) = P(Z<-0.94)= 0.1736

c)

probability = P(72<X<128) = P(-1.75<Z<1.75)= 0.9599-0.0401= 0.9198

d)

P(-1.25<Z<1.25)=0.8944-0.1056=0.7888

2)

mean μ= 40000
standard deviation σ= 4000.0000
for 3th percentile critical value of z= -1.88
therefore corresponding value=mean+z*std deviation= 32480.0